# Online conformal inference for multi-step time series forecasting

### v1

@@ Abstract
[…]
@@ 1 Introduction
Conformal prediction (Vovk et al. 2005) is a powerful and flexible tool for uncertainty quantification, distinguished by its simplicity, generality, and ease of implementation. It constructs valid prediction intervals that achieve nominal coverage without imposing stringent assumptions on the data generating distribution, other than requiring the data to be i.i.d. or, more generally, exchangeable. Its credibility and potential make it widely applicable for quantifying the uncertainty of predictions produced by any black-box machine learning model (Shafer & Vovk 2008, Papadopoulos 2008, Barber et al. 2021) or non-parametric model (Lei & Wasserman 2014).
@@
Three key classes of conformal prediction methods are widely used for constructing distribution-free prediction intervals: split conformal prediction (Vovk et al. 2005), full conformal prediction (Vovk et al. 2005), and jackknife+ (Barber et al. 2021). The split conformal method, which relies on a holdout set, offers computational efficiency but sacrifices some statistical efficiency due to data splitting. In contrast, full conformal prediction avoids data splitting, providing higher accuracy at the cost of increased computational complexity. Jackknife+ strikes a balance between these methods, offering a compromise between statistical precision and computational demands. All three methods guarantee coverage at the target level under the assumption of data exchangeability.
@@
Nevertheless, the data exchangeability assumption is often violated in many applied domains, where challenges such as non-stationarity, distributional drift, temporal and spatial dependencies are prevalent. In response, several extensions to conformal prediction have been proposed to handle non-exchangeable data. Notable examples include methods for handling covariate shift (Tibshirani et al. 2019, Lei & Candès 2021, Yang, Kuchibhotla & Tchetgen Tchetgen 2024), online distribution shift (Gibbs & Candès 2021, Zaffran et al. 2022, Bastani et al. 2022), label shift (Podkopaev & Ramdas 2021), time series data (Chernozhukov et al. 2018, Gibbs & Candès 2021, Xu & Xie 2021, 2023, Zaffran et al. 2022), and spatial prediction (Mao et al. 2024), and methods based on certain distributional assumptions of the data rather than exchangeability (Oliveira et al. 2024, Xu & Xie 2021, 2023). Additionally, some methods propose weighting nonconformity scores (e.g., prediction errors) differently, either using non-data-dependent weights (Barber et al. 2023) or weights based on observed feature values (Tibshirani et al. 2019, Guan 2023, Hore & Barber 2023).
@@
Recently, several attempts have been made to enable conformal prediction on time series data, where exchangeability obviously fails due to inherent temporal dependencies. One line of research has focused on developing conformal-type methods that offer coverage guarantees under certain relaxations of exchangeability. For example, within the full conformal prediction framework, Chernozhukov et al. (2018) and Yu et al. (2022) construct prediction sets for time series by using a group of permutations that are specifically designed to preserve the dependence structure in the data, ensuring validity under weak assumptions on the nonconformity score. In the split conformal prediction framework, Xu & Xie (2021) and Xu & Xie (2023) extend conformal prediction methods under classical nonparametric assumptions to achieve asymptotic valid conditional coverage for time series. Barber et al. (2023) use weighted residual distributions to provide robustness against distribution drift. Additionally, Oliveira et al. (2024) introduce a general framework based on concentration inequalities and data decoupling properties of the data to retain asymptotic coverage guarantees across several dependent data settings.
@@
In a separate strand of research, Gibbs & Candès (2021) develop adaptive conformal inference in an online manner to manage temporal distribution shifts and ensure long-run coverage guarantees. The basic idea is to adapt the miscoverage rate, $\alpha$, based on historical miscoverage frequencies. Follow-up work has refined this idea by introducing time-dependent step sizes to respond to arbitrary distribution shifts, as seen in studies by Bastani et al. (2022), Zaffran et al. (2022), and Gibbs & Candès (2024). However, these methods may produce prediction intervals that are either infinite or null. To address this issue, recent research has proposed a generalized updating process that tracks the quantile of the nonconformity score sequence, as discussed by Bhatnagar et al. (2023), Angelopoulos et al. (2023), and Angelopoulos et al. (2024).
@@
Existing conformal prediction methods for time series primarily focus on single-step forecasting. However, many applications require forecasts for multiple future time steps, not just one. Related research into multi-step time series forecasting is limited and does not account for the temporal dependencies inherent in multi-step forecasts. For example, Stankeviciute et al. (2021) integrate conformal prediction with recurrent neural networks for multi-step forecasting and then apply Bonferroni correction to achieve the desired coverage rate. This approach, however, assumes data independence, which is not often unrealistic for time series data. Yang, Candès & Lei (2024) propose Bellman conformal inference, an extension of adaptive conformal prediction, to control multi-step miscoverage rates simultaneously at each time point $t$ by minimizing a loss function that balances the average interval length across forecast horizons with miscoverage. While this method considers multi-step intervals, it does not account for their temporal dependencies and may be computationally intensive when solving the associated optimisation problems at each time step. Additionally, several extensions to multivariate targets have been explored, see, e.g., Schlembach et al. (2022) and Sun & Yu (2022).
@@
We employ a unified notation to formalize the mathematical representation of conformal prediction for time series data. We consider a general sequential setting in which we observe a time series $\{y_{t}\}_{t\geq 1}$ generated by an unknown data generating process (DGP), which may depend on its own past, along with other exogenous predictors, $\bm{x}_{t}=(x_{1,t},\ldots,x_{p,t})^{\prime}$, and their histories. The distribution of $\{(\bm{x}_{t},y_{t})\}_{t\geq 1}\subseteq\mathbb{R}^{p}\times\mathbb{R}$ is allowed to vary over time, to accommodate non-stationary processes. At each time point $t$, we aim to forecast $H$ steps into the future, providing a prediction set (which is a prediction interval in this setting), $\hat{\mathcal{C}}_{t+h|t}$, for the realisation $y_{t+h}$ for each $h\in[H]$. The $h$-step-ahead forecast uses the previously observed data $\{(\bm{x}_{i},y_{i})\}_{1\leq i\leq t}$ along with the new information of the exogenous predictors $\{\bm{x}_{t+j}\}_{1\leq j\leq h}$. Note that we can generate ex-ante forecasts by using forecasts of the predictors based on information available up to and including time $t$. Alternatively, ex-post forecasts are generated assuming that actual values of the predictors from the forecast period are available. Given a nominal miscoverage rate $\alpha\in(0,1)$ specified by the user, we expect the output $\hat{\mathcal{C}}_{t+h|t}$ to be a valid prediction interval so that $y_{t+h}$ falls within the prediction interval $\hat{\mathcal{C}}_{t+h|t}$ at least $100(1-\alpha)\%$ of the time.
@@
We developed the conformalForecast package for R, available at https://github.com/xqnwang/conformalForecast, to implement the proposed multi-step conformal prediction methods. All the data and code to reproduce the experiments are made available at https://github.com/xqnwang/cpts.
2 Online learning with sequential splits
Let $z_{t}=(\bm{x}_{t},y_{t})$ denote the data point (including the response $y_{t}$ and possibly predictors $\bm{x}_{t}$) at time $t$. Suppose that, at each time $t$, we have a forecasting model $\hat{f}_{t}$ trained using the historical data $z_{1:t}$. Throughout the paper, we assume that the predictors are known into the future. In this way, we perform ex-post forecasting and there is no additional uncertainty introduced from forecasting the exogenous predictors. Using the forecasting model $\hat{f}_{t}$, we are able to produce $H$-step point forecasts, $\{\hat{y}_{t+h|t}\}_{h\in[H]}$, using the future values for the predictors. We define the nonconformity score as the (signed) forecast error
$s_{t+h|t}=\mathcal{S}(z_{1:t},y_{t+h}):=y_{t+h}-\hat{f}_{t}(z_{1:t},\bm{x}_{t+1:h})=y_{t+h}-\hat{y}_{t+h|t}.$
The task is to employ conformal inference to build $H$-step prediction intervals, $\{\hat{\mathcal{C}}_{t+h|t}^{\alpha}(z_{1:t},\bm{x}_{t+1:h})\}_{h\in[H]}$, at the target coverage level $1-\alpha$. For brevity, we will use $\hat{\mathcal{C}}_{t+h|t}^{\alpha}$ to denote the $h$-step-ahead $100(1-\alpha)\%$ prediction interval.
Sequential split. In a time series context, it is inappropriate to perform random splitting, a standard strategy in much of the conformal prediction literature, due to the temporal dependency present in the data. Instead, throughout the conformal prediction methods proposed in this paper, we use a sequential split to preserve the temporal structure. For example, the $t$ available data points, $z_{1:t}$, are sequentially split into two consecutive sets, a proper training set $\mathcal{D}_{\text{tr}}\subset\{1,\ldots,t_{r}\}$ and a calibration set $\mathcal{D}_{\text{cal}}\subset\{t_{r}+1,\ldots,t\}$, where $t_{c}=t-t_{r}\gg H$, $\mathcal{D}_{\text{tr}}\cup\mathcal{D}_{\text{cal}}=\{1,\ldots,t\}$, and $\mathcal{D}_{\text{tr}}\cap\mathcal{D}_{\text{cal}}=\varnothing$. Here, “proper” means that the training set is used exclusively for fitting the model, with no overlap into the calibration set, which is essential for ensuring the validity of coverage in conformal prediction (Papadopoulos et al. 2007). With sequential splitting, multiple $H$-step forecasts and their respective nonconformity scores can be computed on the calibration set.
Online learning. We will adapt the following generic online learning framework for all conformal prediction methods to be discussed in later sections. The entire procedure for the online learning framework with sequential splits is also illustrated in Figure 1. This framework updates prediction intervals as new data points arrive, allowing us to assess their long-run coverage behaviour. It adopts a standard rolling window evaluation strategy, although expanding windows could easily be used instead.
@@ 2 Online learning with sequential splits
[…]
Figure 1: Diagram of the online learning framework with sequential splits. White: unused data; Gray: training data; Black: forecasts in calibration set; Blue: forecasts in test set.
@@ 3 Extending conformal prediction for multi-step forecasting
In this section, we apply the online learning framework outlined in Section 2 to adapt several popular conformal prediction methods in order to allow for multi-step forecasting problems. In time series forecasting, when a model is properly specified and well-trained, the forecasts that minimize the mean squared error (MSE) can be considered optimal in the sense that they achieve the lowest possible expected squared deviation between the forecasts and actual values. One of the key properties of optimal forecast errors, which holds generally in linear models, is that the variance of the forecast error $e_{t+h|t}$ is non-decreasing in $h$ (Tong 1990, Diebold & Lopez 1996, Patton & Timmermann 2007). Therefore, we need to apply a separate conformal prediction procedure for each $h\in[H]$.
@@ 3.1 Online multi-step split conformal prediction
Split conformal prediction (SCP, also called inductive conformal prediction, Papadopoulos et al. 2002, Vovk et al. 2005, Lei et al. 2018), is a holdout method for building prediction intervals using a pre-trained model in regression settings. A key advantage of SCP is its ability to guarantee coverage by assuming data exchangeability. Time series data are inherently nonexchangeable due to their temporal dependence and autocorrelation. Therefore, directly applying SCP to time series data would violate the method’s exchangeability assumption and compromise its coverage guarantee.
Here we introduce online multi-step split conformal prediction (MSCP) as a generalization of SCP to recursively update all $h$-step-ahead prediction intervals over time. Instead of assuming exchangeability, MSCP applies conformal inference in an online fashion, updating prediction intervals as new data points are received. Specifically, for each $h\in[H]$, we consider the following simple online update to construct prediction intervals on the test set:
@@
where $s_{t+h|t}^{y}:=\mathcal{S}(z_{1:t},y)$ denotes the $h$-step-ahead nonconformity score calculated at time $t$ using a hypothesized test observation $y$, $Q_{\tau}(\cdot)$ denotes the $\tau$-quantile of its argument, and $\delta_{a}$ denotes the point mass at $a$.

### v2

@@ Abstract
We consider the problem of constructing distribution-free prediction intervals for multi-step time series forecasting, with a focus on the temporal dependencies inherent in multi-step forecast errors. We establish that the optimal $h$-step-ahead forecast errors exhibit serial correlation up to lag (h1)(h-1) under a general non-stationary autoregressive data generating process. To leverage these properties, we propose the Autocorrelated Multi-step Conformal Prediction (AcMCP) method, which effectively incorporates autocorrelations in multi-step forecast errors, resulting in more statistically efficient prediction intervals. This method guarantees asymptotic marginal coverage for multi-step prediction intervals, though we note that, for finite samples, the coverage error admits an upper bound that increases with the forecasting horizon. Additionally, we extend several easy-to-implement conformal prediction methods, originally designed for single-step forecasting, to accommodate multi-step scenarios. Through empirical evaluations, including simulations and applications to data, we demonstrate that AcMCP achieves coverage that closely aligns with the target within local windows, while providing adaptive prediction intervals that effectively respond to varying conditions.
@@ 1 Introduction
Conformal prediction (Papadopoulos et al. 2002, Vovk et al. 2005) is a simple yet powerful framework for uncertainty quantification. It constructs valid prediction intervals that achieve nominal coverage without imposing stringent assumptions on the data generating distribution, other than requiring the data to be i.i.d. or, more generally, exchangeable. Its credibility and potential make it widely used to quantify uncertainty for predictions produced by black-box machine learning models (Papadopoulos et al. 2007, Papadopoulos 2008, Shafer & Vovk 2008, Barber et al. 2021) or non-parametric models (Lei & Wasserman 2014).
@@
Three widely used classes of conformal prediction methods for constructing distribution-free prediction intervals are split conformal prediction (Vovk et al. 2005), full conformal prediction (Vovk et al. 2005), and jackknife+ (Barber et al. 2021). Split conformal, which relies on a holdout set, offers computational efficiency but sacrifices some statistical efficiency due to data splitting. Full conformal prediction avoids data splitting, providing higher accuracy at the cost of increased computational complexity. Both split and full conformal prediction methods guarantee coverage at the target level under the assumption of data exchangeability. Jackknife+ strikes a balance between these methods, offering a compromise between statistical precision and computational demands. Under assumptions of algorithmic stability, Jackknife+ attains near-nominal coverage, while under the weaker assumption of exchangeability alone, it guarantees coverage of at least $1-2\alpha$ in the worst case. Gupta et al. (2022) further generalize jackknife+ and related out-of-bag conformal methods through a unified framework based on nested families of prediction sets.
@@
Nevertheless, the data exchangeability assumption is often violated in many applied domains, where challenges such as non-stationarity, distributional drift, temporal and spatial dependencies are prevalent. In response, several extensions to conformal prediction have been proposed to handle non-exchangeable data. Notable examples include methods for handling covariate shift (Tibshirani et al. 2019, Yang, Kuchibhotla & Tchetgen Tchetgen 2024), online distribution shift (Gibbs & Candès 2021, 2024, Zaffran et al. 2022, Bastani et al. 2022), time series data (Chernozhukov et al. 2018, Gibbs & Candès 2021, Xu & Xie 2021, 2023, Zaffran et al. 2022), and methods based on certain distributional assumptions of the data rather than exchangeability (Chernozhukov et al. 2018, Oliveira et al. 2024, Xu & Xie 2021, 2023). Additionally, some methods propose weighting nonconformity scores (e.g., prediction errors) differently, either using non-data-dependent weights (Barber et al. 2023) or weights based on observed feature values (Tibshirani et al. 2019, Guan 2023).
@@
Recent work has sought to extend conformal prediction to time series settings, where exchangeability obviously fails due to inherent temporal dependencies. One line of research has focused on developing conformal-type methods that offer coverage guarantees under certain relaxations of exchangeability. For example, within the full conformal prediction framework, Chernozhukov et al. (2018) and Yu et al. (2022) construct prediction sets for time series by using a group of permutations that are specifically designed to preserve the dependence structure in the data, ensuring validity under weak assumptions on the nonconformity score. In the split conformal prediction framework, Xu & Xie (2021) and Xu & Xie (2023) extend conformal prediction methods to time series settings and establish asymptotically valid conditional coverage under the assumption that model errors are stationary and strongly mixing. Barber et al. (2023) use weighted residual distributions to provide robustness against distribution drift. Additionally, Oliveira et al. (2024) introduce a general framework based on concentration inequalities and data decoupling properties of the data to retain asymptotic coverage guarantees across several dependent data settings.
@@
In a separate strand of research, Gibbs & Candès (2021) develop adaptive conformal inference (ACI, denoted as ACP hereafter) in an online manner to manage temporal distribution shifts and ensure long-run coverage guarantees. The basic idea is to adapt the miscoverage rate, $\alpha$, based on historical miscoverage frequencies. However, ACP may yield infinite or empty prediction intervals when the $\alpha$ drifts below 0 or exceeds $1$, respectively. Follow-up work has refined this idea by introducing time-dependent step sizes to respond to arbitrary distribution shifts, as seen in studies by Bastani et al. (2022), Zaffran et al. (2022), and Gibbs & Candès (2024). Recent research has proposed a generalized updating process that tracks the quantile of the nonconformity score sequence, rather than the miscoverage rate, as discussed by Bhatnagar et al. (2023), Angelopoulos et al. (2023), and Angelopoulos et al. (2024).
@@
Existing conformal prediction methods for time series primarily focus on single-step forecasting, even though many applications require reliable uncertainty quantification over multiple future horizons. The literature on multi-step conformal prediction is comparatively limited and, when it does consider multi-step forecasts, it often reduces the problem to $H$ horizon-specific tasks by constructing separate prediction sets for each $y_{t+h}$, $h\in[H]$, without considering how forecasts at different horizons are related. For example, Stankeviciute et al. (2021) integrate conformal prediction with recurrent neural networks for multi-step forecasting and then apply Bonferroni correction to control coverage. This approach, however, assumes data independence, which is often unrealistic for time series. Yang, Candès & Lei (2024) propose Bellman conformal inference to jointly control multi-step miscoverage by minimizing a loss that balances average interval length across horizons and miscoverage, which however ignores temporal dependencies across horizons and can be computationally intensive due to related optimization at each time step. Related extensions to multivariate targets have also been studied; see, for example, Schlembach et al. (2025).
@@
We employ a unified notation to formalize the mathematical representation of conformal prediction for time series data. We consider a general sequential setting in which we observe a time series $\{y_{t}\}_{t\geq 1}$ generated by an unknown data generating process (DGP), which may depend on its own past, along with other exogenous predictors, $\bm{x}_{t}=(x_{1,t},\dots,x_{p,t})^{\prime}$, and their histories. The joint distribution of $\{(\bm{x}_{t},y_{t})\}_{t\geq 1}$, where $\bm{x}_{t}\in\mathbb{R}^{p}$ and $y_{t}\in\mathbb{R}$, is allowed to vary over time, thereby accommodating non-stationary processes. At each time point $t$, we aim to forecast $H$ steps into the future, providing a prediction set (which is a prediction interval in this setting), $\hat{\mathcal{C}}_{t+h|t}$, for the realization $y_{t+h}$ for each $h\in[H]$. The $h$-step-ahead forecast uses the previously observed data $\{(\bm{x}_{i},y_{i})\}_{1\leq i\leq t}$ along with the new information of the exogenous predictors $\{\bm{x}_{t+j}\}_{1\leq j\leq h}$. Note that we can generate ex-ante forecasts by using forecasts of the predictors based on information available up to and including time $t$. Alternatively, ex-post forecasts are generated assuming that actual values of the predictors from the forecast period are available. Given a nominal miscoverage rate $\alpha\in(0,1)$ specified by the user, we expect to construct prediction intervals $\hat{\mathcal{C}}_{t+h|t}$ that achieve long-run coverage guarantees, in the sense that $\lim_{T\rightarrow\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbbm{1}\left\{y_{t+h}\in\hat{\mathcal{C}}_{t+h|t}\right\}\geq 1-\alpha$.
Our goal is to achieve long-run coverage for multi-step univariate time series forecasting. All the proposed methods are grounded in the split conformal prediction framework and an online learning scheme, which are well-suited to the sequential nature of time series data. First, we extend several widely used conformal prediction methods that are originally designed for single-step forecasting to the multi-step setting by constructing horizon-specific prediction intervals. These extensions follow the common practice of treating each horizon independently and, in general, do not provide theoretical long-run coverage guarantees for multi-step prediction intervals, except for the proposed multi-step conformal PID control (MPID) method, which admits such a guarantee. Second, we provide theoretical results showing that, under a general non-stationary autoregressive DGP, the forecast errors of optimal $h$-step-ahead forecasts can be well approximated by a linear combination of at most its lag (h1)(h-1) with respect to forecast horizon. Third, building on the theoretical results, we introduce the autocorrelated multi-step conformal prediction (AcMCP) method, which explicitly accounts for the autocorrelations of multi-step forecast errors at the calibration stage by preserving the dependence structure among nonconformity scores. AcMCP is proven to guarantee asymptotic marginal coverage for multi-step prediction intervals, with the coverage gap converging to zero as the sample size increases, without making any assumptions on data distribution shifts. Our method targets pointwise prediction intervals for each specific forecast horizon, $h\in[H]$. Finally, we illustrate the practical utility of these proposed methods through two simulations and two applications to electricity demand and eating-out expenditure forecasting. The results show that the proposed AcMCP method adapts effectively to changes in the observed data distribution and, among the proposed extensions that achieve coverage close to the target level, generally yields more informative and narrower prediction intervals, particularly for larger forecast horizons.
We developed the conformalForecast package for R to implement the proposed multi-step conformal prediction methods, the package is publicly available on CRAN. All the data and code to reproduce the experiments are made available at https://github.com/xqnwang/cpts.
2 Online learning with sequential splits
Let $z_{t}=(\bm{x}_{t},y_{t})$ denote the data point (including the response $y_{t}$ and possibly predictors $\bm{x}_{t}$) at time $t$. Suppose that, at each time $t$, a forecasting model $\hat{f}_{t}$ is trained on historical data $z_{1:t}$ and updated (re-trained) as new observations arrive. We assume that the predictors are known into the future, corresponding to an ex-post forecasting setting commonly used when predictors are deterministic, policy-driven, or externally specified. By conditioning on future predictors, we focus exclusively on the uncertainty in the response variable $y_{t}$ and avoid introducing additional uncertainty from forecasting the exogenous inputs. Using the forecasting model $\hat{f}_{t}$, we are able to produce $H$-step point forecasts, $\{\hat{y}_{t+h|t}\}_{h\in[H]}$, using the future values for the predictors. We define the nonconformity score as the (signed) forecast error
$s_{t+h|t}=\mathcal{S}(z_{1:t},y_{t+h}):=y_{t+h}-\hat{f}_{t}(z_{1:t},\bm{x}_{(t+1):(t+h)})=y_{t+h}-\hat{y}_{t+h|t}.$
The task is to employ conformal inference to build $H$-step prediction intervals, $\{\hat{\mathcal{C}}_{t+h|t}^{\alpha}(z_{1:t},\bm{x}_{(t+1):(t+h)})\}_{h\in[H]}$, at the target coverage level $1-\alpha$. For brevity, we will use $\hat{\mathcal{C}}_{t+h|t}^{\alpha}$ to denote the $h$-step-ahead $100(1-\alpha)\%$ prediction interval.
Sequential split. In a time series context, it is inappropriate to perform random splitting, a standard strategy in much of the conformal prediction literature, due to the temporal dependency present in the data. Following related work such as Wisniewski et al. (2020) and Zaffran et al. (2022), throughout the conformal prediction methods proposed in this paper, we use a sequential split to preserve the temporal structure. For example, the $t$ available data points, $z_{1:t}$, are sequentially split into two consecutive sets, a proper training set $\mathcal{D}_{\text{tr}}\subset\{1,\dots,t_{r}\}$ and a calibration set $\mathcal{D}_{\text{cal}}\subset\{t_{r}+1,\dots,t\}$, where $t_{c}=t-t_{r}\gg H$, $\mathcal{D}_{\text{tr}}\cup\mathcal{D}_{\text{cal}}=\{1,\dots,t\}$, and $\mathcal{D}_{\text{tr}}\cap\mathcal{D}_{\text{cal}}=\varnothing$. Here, “proper” means that the training set is used exclusively for fitting the model, with no overlap into the calibration set, which is essential for ensuring the validity of coverage in conformal prediction (Papadopoulos et al. 2002, Vovk et al. 2005). With sequential splitting, multiple $H$-step forecasts and their respective nonconformity scores can be computed on the calibration set.
Online learning. We will adapt the following generic online learning framework, which is closely related to Zaffran et al. (2022), for all conformal prediction methods to be discussed in later sections. The entire procedure for the online learning framework with sequential splits is also illustrated in Figure 1. This framework updates prediction intervals as new data points arrive, allowing us to assess their long-run coverage behavior. It adopts a standard rolling window evaluation strategy, although expanding windows could easily be used instead.
@@ 2 Online learning with sequential splits
Figure 1: Online learning framework with sequential splits. White: unused data; Gray: training data; Black: forecasts in calibration set; Blue: forecasts in test set.
@@ 3 Extending conformal prediction for multi-step forecasting
In this section, we apply the online learning framework outlined in Section 2 to extend several popular conformal prediction methods to multi-step forecasting. In time series forecasting, when a model is properly specified and well-trained, the forecasts that minimize the mean squared error can be considered optimal in the sense of achieving the lowest possible expected squared forecast error. A key property of optimal forecast errors, which holds generally for linear models, is that the variance of the forecast error $e_{t+h|t}$ is non-decreasing in $h$ (Diebold & Lopez 1996, Patton & Timmermann 2007). Therefore, a horizon-specific conformal prediction procedure is required for each $h\in[H]$. We note that the extensions of existing approaches introduced in this section do not provide theoretical long-run coverage guarantees, except for the multi-step conformal PID control (MPID) method, which admits such a guarantee.
@@ 3.1 Online multi-step split conformal prediction
Split conformal prediction (SCP, also called inductive conformal prediction, Papadopoulos et al. 2002, Vovk et al. 2005), is a holdout method for building prediction intervals using a pre-trained model on a training set. A key advantage of SCP is its ability to guarantee coverage by assuming data exchangeability. In regression setting, SCP randomly separates the available $n$ data points, $Z_{i}=(X_{i},Y_{i})\in\mathbb{R}^{d}\times\mathbb{R}$, $i=1,\dots,n$, into a proper training set of size $n_{t}$ and a calibration set of size $n_{c}$. Given a regression model $\hat{\mu}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ fitted on the training set and a score function $\mathcal{S}$, nonconformity scores $s_{i}=\mathcal{S}\left(X_{i},Y_{i}\right)$ are computed on the calibration set to measure the nonconformity between the calibration’s response values and the predicted values obtained from the fitted model $\hat{\mu}$. Then SCP computes the prediction interval for the test data $Y_{n+1}$ using
$\hat{\mathcal{C}}_{n+1}^{\alpha}\left(X_{n+1}\right)=\left\{y\in\mathbb{R}:\mathcal{S}\left(X_{n+1},y\right)\leq Q_{1-\alpha}\left(\sum_{i\in\mathcal{D}_{\text{cal}}}\frac{1}{n_{c}+1}\cdot\delta_{s_{i}}+\frac{1}{n_{c}+1}\cdot\delta_{+\infty}\right)\right\},$
where $\mathrm{Q}_{\tau}(\cdot)$ denotes the $\tau$-quantile of its argument, and $\delta_{a}$ denotes the point mass at $a$. Time series data are inherently nonexchangeable due to their temporal dependence and autocorrelation. Therefore, directly applying SCP to time series data would violate the method’s exchangeability assumption and compromise its coverage guarantee.
Here we introduce online multi-step split conformal prediction (MSCP) as a generalization of SCP to recursively update all $h$-step-ahead prediction intervals over time. MSCP applies conformal inference in an online fashion, updating prediction intervals as new data points are received. Specifically, for each $h\in[H]$, we consider the following simple online update to construct prediction intervals on the test set:
@@
where $s_{t+h|t}^{y}:=\mathcal{S}(z_{1:t},y)$ denotes the $h$-step-ahead nonconformity score calculated at time $t$ using a hypothesized test observation $y$.